A racer accelerates from rest at a constant rate of 20 m/s2. ( 8 pts )a. How fast will the racer be going at the end of 6sec?b. How far the first racer traveled during this time ?


1
Expert's answer
2021-11-11T11:59:08-0500

a. "v=v_0+at=0+20\cdot6=120\ (m/s)"


b. "s=at^2/2=20\cdot6^2/2=360\ (m)"